Contribution Margin Calculations. Yard Equipment Company sells leaf blowers for $400 per unit. Variable cost per unit is $80, and fixed costs total $8,000. Find (a) the contribution margin per unit, and (b) the contribution margin ratio.
Answer #1

Contribution margin=Sales-Variable cost  

=(400-80)=$320 per unit

Contribution margin ratio=Contribution margin/Sales

=(320/400)=80%